# Learning to chain-of-thought with Jensen’s evidence lower bound

### v1

@@ 8.2 Ablation study
Multi-sample ablation on sample size $n$.
We ablate on the number of sample $n$ used for constructing per gradient update. In theory, as $n$ increases, the multi-sample lower bound becomes tighter and asymptotically approaches the marginal likelihood objective (which is equivalent to the RL objective). We vary the sample size $n\in\{1,2,4,8\}$ and compare the performance. Figure 3 shows that as $n$ increases, the algorithm becomes more KL-efficient: with a fixed budget on KL, the model obtains better performance. Intriguingly, we also observe a training performance akin to reward over-optimization (Gao et al., 2023) - as the optimization progresses, the training reward drops slightly (for blue curve). We can interpret this as the result of the fact that lower bound algorithms do not optimize for the same indicator matching function as the reward we monitor.
Regularization ablation.
@@
Figure 5: Ablation of model size (8B vs. 70B). We find that the multi-sample lower bound algorithm is fairly competitive against the online RL algorithm in the 70B scale. Both algorithm traces out a similar KL-performance trade-off, with multi-sample lower bound algorithm obtaining a slightly better performance given a similar compute budget as online RL.
Scaling up model size.
@@
Since the multi-sample lower bound algorithm appears more competitive, we compare it against the online RL in the 70B case. Figure 5 shows that the lower bound algorithm obtains competitive performance against online RL in terms of the KL-performance trade-off. With roughly the same amount of compute budget, we find that the lower bound algorithm seems to drift further from the reference policy, hence extending the trade-off curve to a performance of $70\%$ test set accuracy, which outperforms online RL modestly.
9 Conclusion and limitations
While the theoretical perspective that views chain-of-thought as latent variable is not new, we propose a much practically simpler approach to maximizing the evidence lower bound. The lower bound technique draws intriguing connections between SFT, RL, and sheds light on how literature on probabilistic inference produces algorithmic insights to optimizing chain-of-thought.
Possible directions for future research include studying how such algorithms can be applied to domains where good reward is difficult to come by, and how they scale with data quality compared to RL, how the approach might generalize to hidden chain-of-thought (Jaech et al., 2024; Xiang et al., 2025), and how the algorithm might combine with a reward function or verifier that generates ground truth answer in an iterative fashion.
APPENDICES: Learning to chain-of-thought with Jensen’s evidence lower bound
@@ Appendix A A review of the graphical model perspective
Probabilistic inference with a learnable prior.
@@
$\log p_{\theta}(o)\geq\underbrace{\mathbb{E}_{z\sim q_{\theta}(\cdot|o)}\left[\log p_{\theta}(z|o)+\log\frac{q_{\phi}(z|o)}{p_{\theta}(z)}\right]}_{\mathcal{L}_{\theta,\phi}(o)}.$
@@
Chain-of-thought with full ELBO.
@@
$\log\pi_{\theta}(a^{*}|x)\geq\underbrace{\mathbb{E}_{c\sim q_{\theta}(\cdot|x,a^{*})}\left[\log\pi_{\theta}(c|x,a^{*})+\log\frac{q_{\phi}(c|x,a^{*})}{\pi_{\theta}(c|x)}\right]}_{\mathcal{L}_{\theta,\phi}(x,a^{*})}.$
Chain-of-thought with Jensen’s lower bound.
@@
Jensen’s lower bound with regularization.
@@ Appendix B Hyper-parameters and experimental settings
We train on the MATH training set with $7500$ examples and evaluate on the test set with $2500$ examples. A supervised fine-tuning on the training set is conducted to warm up the RL training, hence the gap between training and test set accuracy.
@@
For both training and evaluation, we provide system instructions that ask the model to generate a response with step-by-step solution, followed by a final conclusion phrased as the final answer is followed by the answer predicted by the model. This is consistent with the prompt structure discussed for Llama models [Dubey et al., 2024, Yue et al., 2024].
All experiments are conducted with an entropy regularization coefficient $\beta>0$ which we have ablated in the main paper.
B.1 Lower bound algorithm details
We add an additional RL loss with the reward as $r_{\text{reg}}(x,y)=-10$ if $y$ does not follow the formatting requirement (that the identifier phrase the final answer is is in $y$) and zero otherwise. We find that this generally helps stabilize the lower bound training process. This is especially useful for small models where under temperature sampling, it can often not follow instructions strictly.
The lower bound loss is applied only to generations that abide by the formatting requirement, i.e., with the final answer is found inside the full generation. Otherwise, the loss is masked out. After generating a response $y$ from the model, we can parse out its chain-of-thought component $y=(c,a)$ and replace its generated answer with the ground truth. This produces another sequence $y^{*}=(c,a^{*})$. Running a forward pass on such a sequence produces quantities such as $\log\pi_{\theta}(c|x)$ and $\log\pi_{\theta}(a^{*}|c,x)$ required for the algorithm.
Advantage post-processing.
Both the baseline RL and the lower bound algorithm applies advantage post-processing, following common practice [Schulman et al., 2017]. For example, in the multi-sample lower bound algorithm, the advantage for the $i$-th generation is
$A_{i}=\log\left(\frac{1}{n}\sum_{j=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{j})\right)-\tilde{v}_{i}.$
A further normalization is applied to the advantage $\tilde{A}_{i}=\text{clip}(A_{i}/\text{std}\left(A\right),-1,1)$ such that the outcome $\tilde{A}_{i}$ is applied in the actual update. For the baseline RL algorithm, a similar procedure applies. Put together, the fully implemented multi-sample lower bound update is
@@ B.1 Lower bound algorithm details
$\frac{1}{n}\sum_{i=1}^{n}\left(\left(\tilde{A}_{i}+\tilde{A}_{i}^{\text{(ref)}}\right)\nabla_{\theta}\log\pi_{\theta}(a_{i}|x,c_{i})\right)+\nabla_{\theta}\log\frac{1}{n}\sum_{i=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{i})-\beta\nabla_{\theta}\mathbb{KL}\left(\pi_{\theta}(\cdot|x),\pi_{\text{ref}}(\cdot|x)\right),$
where $\tilde{A}_{i}^{\text{ref}}=\text{clip}\left(\left(r_{\text{ref}}(x,y_{i})-\bar{r}_{\text{ref}}\right)/\text{std}(r_{\text{reg}}),-1,1\right)$ is the normalized advantage for the formatting penalty. The normalization makes it such that the ultimate update optimizes for a lower bound more akin to the weighted lower bound [Higgins et al., 2017] though the underlying algorithmic motivations differ.
Appendix C Variance-reduced RL gradient estimate
We provide more discussion on the variance-reduced RL gradient estimate.
C.1 Proof of variance reduction

### v2

@@ 8.2 Ablation study
Multi-sample ablation on sample size $n$
We ablate on the number of sample $n$ used for constructing per gradient update. In theory, as $n$ increases, the multi-sample lower bound becomes tighter and asymptotically approaches the marginal likelihood objective (which is equivalent to the RL objective). We vary the sample size $n\in\{1,2,4,8\}$ and compare the performance. Figure 4 shows that as $n$ increases, the algorithm becomes more KL-efficient: with a fixed budget on KL, the model obtains better performance. Intriguingly, we also observe a training performance akin to reward over-optimization (Gao et al., 2023) - as the optimization progresses, the training reward drops slightly (for blue curve). We can interpret this as the result of the fact that JEPO does not optimize for the same indicator matching function as the reward we monitor.
Regularization ablation
@@
Figure 6: Ablation of model size (8B vs. 70B). We find that the multi-sample JEPO is fairly competitive against the online RL algorithm in the 70B scale. Both algorithm traces out a similar KL-performance trade-off, with multi-sample JEPO obtaining a slightly better performance given a similar compute budget as online RL.
Figure 7: Evaluation comparison of training 70B models on semi-verifiable Numina dataset. We show evaluation results during the course of training. Left plot shows the combined accuracy on the unverifiable subset (about 40%) of the test set; middle plot shows the combined accuracy on the full test set; right plot shows the Sympy score on the full test set. While JEPO progresses more slowly on the Sympy scores compared to online RL, it gains on the combined accuracy; the combined algorithm seems to achieve the best of both worlds.
Scaling up model size
@@
Since the multi-sample JEPO appears more competitive, we compare it against the online RL in the 70B case. Figure 6 shows that the JEPO obtains competitive performance against online RL in terms of the KL-performance trade-off. With roughly the same amount of compute budget, we find that the JEPO seems to drift further from the reference policy, hence extending the trade-off curve to a performance of $70\%$ test set accuracy, which outperforms online RL modestly.
Supervised loss
We find that a low value of $\beta_{\text{sup}}$ generally works better for the JEPO algorithms. The speculation is that when $\beta_{\text{sup}}$ is large, the supervised loss encourages the policy to place weights on the ground truth $a^{*}$ despite that the chain-of-thought $c$ has low quality. This leads to overfitting the training set, in a more severe way than online RL. This is because the JEPO supervised learning loss incentivizes the model to directly memorize $a^{*}$ given any context $(x,c)$, by maximizing the likelihood $\log\pi_{\theta}(a^{*}|x,c)$.
Interestingly, we will show that with long-form data, large values of $\beta_{\text{sup}}$ work generally better, since the risk of overfitting is less severe.
9 Experiments with semi-verifiable data
We now consider semi-verifiable data where a good proportion of the dataset contains answers which are not easily verifiable. We focus on a post-processed Numina dataset (LI et al., 2024) where prompts are mathematical questions and ground truth answers are partly verifiable. For instance, one example of the ground truth answer is the whole expression: $\forall x\in\mathbb{R},x^{2}+(a-1)x+1\geq 0$. Given a model generated answer, it is hard to verify whether it is equivalent to the above expression without case-specific parsing; often time, such parsing results in false negatives. See Appendix B for details on how we post-process the dataset and data splits.
RL baseline and reward
For the RL baseline, we apply the Sympy reward as introduced in the previous section. Because the dataset contains answers which are hard to verify, the reward is effectively only applicable to a subset of the data. The default training set contains about 22k examples. We estimated at least 40% of such examples cannot be verified by the automatic scorer. We consider online RL with such reward as a baseline, as it has access to a highly specialized verifiable reward but only applicable to a subset of the data.
Combining JEPO and RL baseline
We also compare with an algorithm that combines the loss function of JEPO and RL baseline with the Sympy reward. When we sample $n$ generations from a single prompt, and if none of the generation obtains a positive score (note this does not mean that the example is necessarily unverifiable), we apply the JEPO loss; otherwise, we apply the baseline RL loss. This allows for a dynamic combination of two losses, and still leverages the whole dataset.
Evaluation
As with the trainging set, the held-out test set contains both verifiable and unverifiable examples, which we evaluate in two ways: (1) Sympy reward $r_{\text{sympy}}(a,a^{*})\in\{0,1\}$, which generally underestimates the true accuracy when ground truth is semi-verifiable; (2) Sympy combined with LLM-as-judge $r_{\text{combined}}(a,a^{*})$, which combines two sources of scores
$r_{\text{combined}}(a,a^{*})\coloneqq r_{\text{sympy}}(a,a^{*})+r_{\text{llm}}(a,a^{*})\mathds{1}_{\{r_{\text{sympy}}(a,a^{*})=0\}}.$
The LLM-as-judge score $r_{\text{llm}}(a,a^{*})$ is also binary: it is based on a 5-time majority voted decision of a prompted 70B instruction-tuned model (Dubey et al., 2024). Though imperfect, we observe that LLM-as-judge reasonably mitigates some false negatives caused by rigid Sympy scoring. Importantly, we reiterate that we do not train on such combined scores - they are used for evaluations only.
9.1 Comparison on Numina
Unless otherwise stated, we will experiment with the multi-sample algorithm given its performance gains in Section 8. Below, Figure 7 shows the evaluation performance comparing the RL baseline, JEPO and their combined algorithm. Since the Numina dataset is more challenging, we experiment throughout with 70B models.
Sympy scoring evaluation
Figure 7 right plot shows the evaluation accuracy using the Sympy score. Overall, all algorithms make steady progress as the training progresses. However, since online RL baseline trains with the same reward signal, it achieves slight acceleration compared to JEPO. The combined algorithm achieves a similar rate of progress with the Sympy scores on the test set.
Due to the abundance of symbolic expressions as ground truth answers in the Numina dataset, here the Sympy reward is a much more specialized scoring method than e.g., string match compared to the MATH case. This partly explains why the online RL baseline is quite competitive, as it also trains on the very same signal. Note this experimental result corroborates the trade-off discussion in Section 5.
Combined scoring evaluation
Figure 7 left plot and middle plot shows the combined accuracy which alleviates some false negatives due to the Sympy scoring. As seen from the overall metrics, the accuracy increases by about 25% compared to the Sympy scores. The left plot shows the performance on the unverifiable test subset (40% of the test set) while the middle plot shows the full set. We observe that both JEPO and the combined algorithm achieves faster rate of progress and asymptotes to slightly better performance than the online RL baseline with this combined metric, especially on the unverifiable subset. Interestingly, note that by training on verifiable rewards, online RL can also make progress on the unverifiable test set.
Though the Sympy scoring is quite specialized, it is only applicable to a subset of the full Numina training set. Meanwhile, JEPO can leverage the full dataset, despite with less specialized signals. The combined algorithm seems to achieve the best of both worlds.
Figure 8: Test set proxy NLL evaluation for training on the numina dataset. We evaluate the proxy NLL of the trained models on the numina test set, approximated with $n=4$ samples lower bound defined in Eqn (9). Both JEPO and the combined algorithm sees improvement in the NLL (lower the better), while online RL does not improve on test set NLL. This hints at different solutions found by the online RL and JEPO algorithms, despite similar improvement trend in the sampling based evaluations.
9.2 Ablation study
We discuss a few additional ablations on the Numina dataset.
Test set negative likelihood: lower the better
We further evaluate the proxy negative log likelihood (NLL) that the trained model produces on test set, computed via the $n$-sample lower bound
[…]
where the expectation is under $(c_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x),(x,a^{*})\sim\mathcal{D}_{\text{test}}$, following common practice (Burda et al., 2015; Ruan et al., 2025). Figure 8 shows such proxy NLL during training, where we see a different pattern for the online RL baseline and JEPO. For JEPO, the proxy NLL decreases over time. We expect such a result because JEPO optimizes for the same objective on the training set, and before overfitting, we expect improvement on the test set.
Meanwhile, maybe surprisingly, online RL does not make progress on the test set NLL. The combined algorithm is in between the two extremes. There are good reasons for online RL not to make progress on test set NLL. Particularly, for each ground truth answer in the dataset $a^{*}$, the Sympy scorer defines a sizable collection of correct answers $\mathcal{A}=\{a:r_{\text{sympy}}(a,a^{*})=1\}$ whose aggregate probability $\pi_{\theta}(\mathcal{A}|x)$ increases under online RL (evidenced by test set accuracy improvement in Figure 7 right plot). In other words, online RL might not improve the proxy NLL of a particular $a^{*}$ (defined through the dataset) inside $\mathcal{A}$.
The above observation implies that the policy found by online RL and JEPO can produce different answers to the same question. It is also suggestive of how reward-based RL post-training can change the calibration behavior of likelihood-based models (Achiam et al., 2023).
Comparison with SFT baseline on golden chain-of-thought
To assess another option to improve semi-verifiable performance, we carry out another comparison against a SFT baseline, which trains on the golden chain-of-thought found in the source dataset (LI et al., 2024). We observe performance improvements across evaluation metrics as well, though generally under-performing RL. See Appendix C for full results.
10 Experiments with unverifiable data
Finally, we experiment with unverifiable data, where the full dataset has long-form ground truth and whose equivalence against another solution cannot be easily verified with hard-coded programs. We consider a post-processed Numina-proof, extracted from the original Numina dataset. The proof often contains multiple sentences or paragraphs, without a final short-form answer as in MATH or the verifiable subset of Numina.
Figure 9: Test set proxy NLL evaluation for training on the unverifiable Numina-proof dataset. For JEPO outperforms the SFT baseline at the same data budget (measured in epochs) and achieves asymptotically better test performance.
Baselines
Since the ground truth is long-form and cannot be verified easily, we do not have a RL baseline with verifiable reward. Instead, SFT on the raw dataset $(x,a^{*})$ is a reasonable baseline. Through a few ablations, we also compare with methods akin to VR-CLI (Gurung and Lapata, 2025), which corresponds to the REINFORCE part of the single-sample lower bound gradient in Eqn (6).
Evaluation
We evaluate NLL on the test set, akin to the ablations in Section 9. We do not carry out sampling based evaluations as long-form answers are hard to assess even for frontier models (Petrov et al., 2025).
10.1 Comparison on Numina-proof
As main experiments, we compare JEPO with SFT. Note that we always started with instruction-tuned models (Dubey et al., 2024) and the SFT baseline can be understood as a continued SFT. We show the curve after an initial transient phase where the test set NLL drops significantly for all runs, which can be attributed to that the modes learn to format answer correctly.
Figure 9 shows the test set NLL comparison between SFT and JEPO, with both 8B and 70B models. At both scales, JEPO outperforms SFT with test set NLL at the same training data epoch. Meanwhile, JEPO also achieves asymptotically better NLL than SFT.
10.2 Ablation study
Figure 10: Comparison of different baselines on numina-proof test set NLL, across various algorithmic variants, with the 70B model. We observe that the supervised component of the JEPO loss plays a key role at learning efficiency and achieving good asymptotic performance.
To understand the role of each loss component, we carry out a few additional comparisons. Recall that JEPO update contains two parts: a REINFORCE component, whose single-sample variant is akin to VR-CLI (Gurung and Lapata, 2025); and a supervised loss component. We compare with a variant where the supervised loss is down-weighted ($\beta_{\text{sup}}=0.01$) and another where it is removed ($\beta_{\text{sup}}=0$).
Figure 10 shows the comparison on the test set NLL. We see that by downweighting the supervised loss, JEPO makes much less progress on the test NLL given the same training epochs. Specifically, when the supervised loss is removed ($\beta_{\text{sup}}=0$), test NLL also seems to plateau at a worse level. Interestingly, this contrasts the observation in MATH experiments (Section 8) where small values of $\beta_{\text{sup}}$ work better. We speculate that the one difference is that nature of the chain-of-thoughts differs: for MATH or general short-form mathematical QA, the chain-of-thought details solution steps and a final answer can be readily inferred. For long-form data, the chain-of-thought tends to be a high-level outline, and it still takes extra effort to produce the full answer (e.g., proof). For the latter case, the supervised learning loss is useful.
11 Conclusion, discussions and limitations
We propose JEPO, a generic training paradigm scaling RL to unverifiable data, without the need for external verifiable rewards. We focus on the case where the reward is computed by matching a model generated solution against a dataset ground truth. We heavily draw on the probabilistic inference formulation that views chain-of-thought as latent variable. Bypassing the modeling complexity required for full ELBO, we propose to multi-sample Jensen’s evidence lower bound for scalable training. We show competitive performance on a wide array of datasets, ranging from verifiable data like short-form math problems to unverifiable data like proof.
Possible directions for future research include studying the impact that various loss components (e.g., the REINFORCE and the supervised loss) have on overfitting; more organic ways to combine verifiable rewards and JEPO; and ways to scale JEPO in the form of meta-thought (Jaech et al., 2024; Xiang et al., 2025) or to pre-training (Ruan et al., 2025).
@@ Appendix A A review of the graphical model perspective
Probabilistic inference with a learnable prior
@@
$\log p_{\theta}(o)\geq\underbrace{\mathbb{E}_{z\sim q_{\theta}(\cdot|o)}\left[\log p_{\theta}(o|z)+\log\frac{q_{\phi}(z|o)}{p_{\theta}(z)}\right]}_{\mathcal{L}_{\theta,\phi}(o)}.$
@@
Chain-of-thought with full ELBO
@@
$\log\pi_{\theta}(a^{*}|x)\geq\underbrace{\mathbb{E}_{c\sim q_{\theta}(\cdot|x,a^{*})}\left[\log\pi_{\theta}(a^{*}|x,c)-\log\frac{q_{\phi}(c|x,a^{*})}{\pi_{\theta}(c|x)}\right]}_{\mathcal{L}_{\theta,\phi}(x,a^{*})}.$
Chain-of-thought with Jensen’s lower bound
@@
Jensen’s lower bound with regularization
@@ Appendix B Hyper-parameters and experimental settings
We experimented with the Llama 3.1 model of size 8B and 70B. All experiments are conducted with identical hyper-parameter settings: we always applied a batch size of $B=128$ prompts per 8B update and $B=64$ per 70B update, and sampled $n=4$ distinct generations per prompt. We found these hyper-parameters so that the model fits the GPU group memory as much as possible.
@@
For the verifiable experiments, a supervised fine-tuning on the training set was conducted to warm up the RL training, hence the beginning gap between training and test set accuracy. For the semi-verifiable and unverifiable experiments, we directly apply JEPO to the released checkpoints.
All updates were carried out with the Adam optimizer [Kingma and Ba, 2014] with learning rate $4\cdot 10^{-7}$. We found this learning rate by starting from a smaller value and increased the learning rate $2x$ at each iteration, to see if the training speeds up without hurting performance. We expect the results to be somehow robust to slight changes in learning rates.
@@
Throughout all experiments (verifiable, semi-verifiable and unverifiable), for both training and evaluation, we provide system instructions that ask the model to generate a response with step-by-step solution, followed by a final conclusion phrased as the final answer is followed by the answer predicted by the model. This is consistent with the prompt structure discussed for Llama models [Dubey et al., 2024, Yue et al., 2024].
B.1 Training and evaluation dataset
For the verifiable experiments on MATH, we train on the MATH training set with $7500$ examples and evaluate on the test set with $2500$ examples [Hendrycks et al., 2021]. For the semi-verifiable experiments, we train with the post-processed Numina dataset [LI et al., 2024], where we split the post-processed 22k examples into a training set (90%) and test set for evaluation. For unverifiable experiments, we extract the proof specific subset from the full Numina dataset, and split training and test set the same way as before.
B.2 Numina dataset post-processing
We use unverifiable proofs data from Numina 1.5 [LI et al., 2024] for our experiments. We clean and filter the questions and their corresponding solutions using some simple regex heuristics. For example, we replace leading blanks, markdown headings like ##, prefixes like “Problem:” and “Solution”, letter-digit combinations like “A1” / “G5” / “ROU”, and trailing dots and blanks. After cleaning, we have 58088 proofs from the Numina dataset.
B.3 Online RL baseline
The online RL baseline is implemented akin to prior work such as RLOO [Ahmadian et al., 2024], which can be understood as an on-policy special case of GRPO [Shao et al., 2024]. Specifically, given the verifiable reward, $r_{i}$, the advantage is computed with standard post-processing $A_{i}=\text{clip}\left((r_{i}-\bar{r}_{-i})/\text{std}(r_{i}),-1,1\right)$ where $\bar{r}_{-i}$ is the leave-one-out control variate. In sum, The update is
@@ B.3 Online RL baseline
$\frac{1}{n}\sum_{i=1}^{n}\left(A_{i}\cdot\nabla_{\theta}\log\pi_{\theta}(a_{i}|x,c_{i})\right)+A_{i}\cdot\nabla_{\theta}\log\pi_{\theta}(a_{i}|x,c_{i})-\beta\nabla_{\theta}\mathbb{KL}\left(\pi_{\theta}(\cdot|x),\pi_{\text{ref}}(\cdot|x)\right),$
where we intentionally separate the update to the chain-of-thought $c_{i}$ and sampled answer $a_{i}$. Juxtaposing the above update with the JEPO update, we note that the supervised loss $\nabla\log\pi_{\theta}(a^{*}|x,c_{i})$ in JEPO echos the answer update $A_{i}\cdot\nabla_{\theta}\log\pi_{\theta}(a_{i}|x,c_{i})$ in the RL algorithm. Formally, we show that the connection between the supervised loss and a variance reduced variant to the online RL update (Section 5), see also Figure 1 for a summary of high-level comparison.
Appendix C Additional ablations on semi-verifiable data
Figure 11: Additional comparison against a SFT baseline which trains on golden chain-of-thought data from the numina dataset. We show that the SFT baseline also improves upon various metrics, despite generally underperforming RL algorithms.
We show the comparison against a SFT baseline on golden chain-of-thought dataset in Figure 11. A few observations are in order: (1) SFT generally is not as good as the RL jobs, but it improves over time as we train more; (2) There is an initial drop in performance, which can be explained by the fact that the golden chain-of-thought does not conform to the familiar step-by-step that the starting model has been post-trained with [Dubey et al., 2024]. Through SFT, the model needs to unlearn the step-by-step format and learns the more freeform hybrid format in the golden chain-of-thought data; (3) Asymptotically, SFT performs lower than RL runs.
Appendix D Proof of variance reduction for variance-reduced RL gradient estimate